OpenStudy (anonymous):

What is the simplified form of (1/x+1/y)/ (1/x-1/y) ?

OpenStudy (anonymous):

\[\large\frac{\frac{1}{x}+\frac{1}{y}}{\frac{1}{x}-\frac{1}{y}}\] The LCD is xy, so we multiply numerator and denominator by xy \[\frac{y+x}{y-x}\]

OpenStudy (anonymous):

Thank Youu (: